X2+5X+6=0 by complating square
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

X^2+5X+6=0
X^2+5X = -6
X^2+5X+25/4 = -6+25/4 (complete square by adding 25/4 to both sides)
(X+5/2)^2 = -24/4 + 25/4
(X+5/2)^2 = (-24+25)/4
(X+5/2)^2 = 1/4
X+5/2 = 1/2
X = 1/2 - 5/2
X = -4/2
X = -2


Answer by stanbon(75887)

X^2+5X+6=0 by completing the square
-----
x^2 + 5x + (5/2)^2 = -6+(5/2)^2
-----
(x+(5/2))^2 = -24/4 + 25/4
----
x+(5/2)^2 = 1/4
---
x + (5/2) = 1/2 or x+(5/2) = -1/2
---
x = -2 or x = -3
====================
Cheers,
Stan H.
===============